Hi,
I've been having trouble with this word problem and I honestly don't know where to start. Your help would be greatly a appreciated.
Next-door neighbors Bob and Jim use hoses from both houses to fill Bob's swimming pool. They know it takes 18 hours using both hoses. They also know that Bob's hose, used alone, takes 20% less time than Jim's hose alone. How much time is required to fill the pool by each hose alone?

Found 2 solutions by rfer, edjones:
Answer by rfer(16322)
Answer by edjones(8007)

Let b=time it takes bob's hose and j=jim's hose's time
b=.8j
1/b + 1/j = 1/18
1/.8j + 1/j = 1/18
18+14.4=0.8j Multiply each side by LCM=14.4 to eliminate fractions.
.8j=32.4
j=40.5 hrs
b=32.4 hrs
.
Ed